OpenStudy (anonymous):

In 1990, when the consumer price index (CPI) was 130.7, Deena purchased a house for $98,700. Assuming that the price of houses increased at the same rate as the CPI from 1980 to 1990, approximately how much would the house have cost in 1980, when the CPI was 82.4?

OpenStudy (ranga):

This is a proportions problem. Let x be the cost of the house in 1980. Year 1980 1990 Cost of House x $98,700 CPI 82.4 130.7 x / 82.4 = 98700 / 130.7 x = ?